OpenStudy (anonymous):

What are the zeros of x^3+8=0? (hint: factor and solve)

OpenStudy (aravindg):

\[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

factor as the sum of two cubes \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

in your case you get \[(x+2)(x^2-2x+4)=0\]

OpenStudy (anonymous):

first factor gives the real zero \(x=-2\) second factor gives two complex zeros via the quadratic formula

OpenStudy (anonymous):

if you are not looking for complex zeros, then it would be a lot easier to write \[x^3=-8\] and get \(x=-2\) by inspection

OpenStudy (anonymous):

how do you use the quadratic formula in this?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

oh i understand, thank you :]